Once upon a time, the treasurer of “Mighty Corporation” (MCO) decided to issue a bond (hereafter: The bond) The bond would have a 20-year life and promised that the holder of the bond would receive:

  • 20 annual payments of $80,
  • Along with the 20th payment, MCO promised to return the principal of $1,000. In other words, the final payments received at maturity would be ($1000+$80) = $1080.
  • On the day it is issued, the Coupon Yield is 80/1000= 8%
  • On the day it is issued, the Yield-to-Maturity (YTM) is 8%

The bond is bought by an insurance company. One-year passes, the bond has made its first coupon payment, and the interest rate/yield-to-maturity on the bond falls to 6%.

A) What is the Coupon Yield on the bond?

B) The bond now has 19 years left. Calculate the Macaulay duration of the bond.
Answer #1

A

                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =19
Bond Price =∑ [(8*1000/100)/(1 + 6/100)^k]     +   1000/(1 + 6/100)^19
                   k=1
Bond Price = 1223.16
Using Calculator: press buttons "2ND"+"FV" then assign
PMT = Par value * coupon %=1000*8/(100)
I/Y =6
N =19
FV =1000
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(6/(100),19,-8*1000/(100),-1000,)
current yield = coupon rate*par value/current price
Current yield%=(8/100)*1000/1223.16
Current yield% = 6.54

B

php3FGfSr.png

Period Cash Flow Discounting factor PV Cash Flow Duration Calc
0 ($1,223.16) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period
1                 80.00                                                             1.06                    75.47                  75.47
2                 80.00                                                             1.12                    71.20                142.40
3                 80.00                                                             1.19                    67.17                201.51
4                 80.00                                                             1.26                    63.37                253.47
5                 80.00                                                             1.34                    59.78                298.90
6                 80.00                                                             1.42                    56.40                338.38
7                 80.00                                                             1.50                    53.20                372.43
8                 80.00                                                             1.59                    50.19                401.54
9                 80.00                                                             1.69                    47.35                426.17
10                 80.00                                                             1.79                    44.67                446.72
11                 80.00                                                             1.90                    42.14                463.57
12                 80.00                                                             2.01                    39.76                477.09
13                 80.00                                                             2.13                    37.51                487.59
14                 80.00                                                             2.26                    35.38                495.38
15                 80.00                                                             2.40                    33.38                500.72
16                 80.00                                                             2.54                    31.49                503.87
17                 80.00                                                             2.69                    29.71                505.06
18                 80.00                                                             2.85                    28.03                504.50
19           1,080.00                                                             3.03                  356.95              6,782.13
      Total            13,676.89
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=13676.89/(1223.16*1)
=11.181602